Paul and Mary are in the 37% ax backet tor ordinary n come and the 20% brac t tor capital gains ignore the 3.8% additional tax on vest en income t higher income taxpayers. They have owned several blocks of stock for many years. They are considering the sale of thwo blocks of stock. The sale of one block would produce a gain of $12,000. The sale of the other would produce a loss of $20,000. For purposes of this problem, ignore any restrictions on deductions or credits based on AGI. They have no other gains or losses this year Read the requirements Requirement a. How much tax wil they save if they sell the block of stock that produces a loss? Paul and Mary will save Requirement b. How much additional tax will they pay if they sell the block of stock that produces a gain? Paul and Mary will pay an additional 1,110
Answer #1

a) They can only recognize 3000 loss, so their tax savings would be 3000*37% = 1110

b) The additional tax they will pay if they sell the block of stock that produces a gain = 12000*20% = 2400

thank you